An initial deposit of $3000 is made in a savings account for which the interest is compounded continuously. The balance will double in seven years. What is the annual rate of interest for this account?
Answer by stanbon(75887)

Formula for continuously compounded investment: A=Pe^(rt)
If deposit doubles in 7 years:
2P=Pe^(7r)
2=e^(7r)
ln(2)-7r
r=ln(2)/7=9.9%
Cheers,
Stan H.